OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. Answer x intercept = (21, 0); y intercept = (0, 3) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (3, 0); y intercept = (0, 21) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

Well first to find it, you want to get from the Standard version of the equation (Ax+By=C) which is where you have it now, to y-intercept version (y=mx+b) where b is the y intercept. x+7y=-21 -x -x 7y=-x-21 /7 /7 /7 y=-1/7x-3 so the y intercept is -3. That narrows your answers down right to the last option, x intercept=(-21,0) y intercept= (0,-3)